a concrete swimming pool is a rectangular prism with the following dimensions: length of 6 metres, width of 4 metres and depth of 1.3 metres.
a) what surface area of tiles is needed to line the inside of the pool (remember there is no top on the pool. give answer in m2.)
b)how much will it cost to line the pool if it costs $40 per m2??

Answer by macston(5194)

L=length; W=width; d=depth
.
The surface area (SA)of the pool can be given by:
SA=(perimeter*depth) + (length*width)
SA=((2(L+W))*d)+(LW)
SA=((2(6+4))*1.3)+(6*4)
SA=26+24=50m^2
ANSWER a: Area of tiles needed is 50 meters squared.
.
($40/m^2)(50m^2)=$2000
ANSWER b: It will cost $2000 to line the pool.